Find the appropriate distribution for this question. There 20 multiple choice questions in a Statistic class quiz. Each question has 4 possible answers then only 1 of them has correct answer. Find the probability of having 3 or less correct answer if a student has to answer every question in random. Give an explanation on the result.


1
Expert's answer
2022-01-13T13:05:56-0500

The probability of guessing the correct answer out of 4 answers is ¼ or 0.25. That makes the probability of guessing the wrong answer ¾ or 0.75. Assuming the 20 questions are independent of each other, this problem is represented by a binomial probability distribution where n = 20, p = ¼, q = ¾.

for binomial probability distribution:

"P(x=k)=C^k_n p^k q^{n-k}"

Hence, the probability of getting 3 or less correct answer:

"P(x\le 3)=P(0)+P(1)+P(2)+P(3)"

where

"P(0)=0.75^{20}=0.0032"

"P(1)=20\cdot0.25\cdot 0.75^{19}=0.0211"

"P(2)=190\cdot0.25^2\cdot 0.75^{18}=0.0167"

"P(3)=\frac{20!}{17!3!}\cdot 0.25^3 \cdot 0.75^{17}=0.1339"


"P(x\le 3)=0.0032+0.0211+0.0167+0.1339=0.1749"